• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name bogdanw

    Log entry time 10:50:28 on December22,2003

    Entry number 115031


    keyword=A1/A2 few channels were fixed

    A2 PMT -a2.tdc[8] - noise ~ 30 kHz -> after pulses due He - need to be replace before experiment.
    A2 PMT -a2.adc[10] in analyzer- HV base failed - replaced, looks good
    A2 PMT -a2.tdc[16] - was empty - signal cable on the discriminator wasn't locked
    A1 PMT -a1.adc[5]-[6] - HV cables on the PMT were misplaced - fixed
    A1 PMT -a1.adc[7]) big current in base - base replaced, current is low

    Present HV set is in the picture

    FIGURE 1